Find the absolute min/max of f(,y) = x2 + y2 +r+y + 4 on the domain D:((x,y):|x|\leq 1,y\leq 1)
Answer #1

fla,y) = x² + y² + x²y +4 of 2x + 2xy 2 - 2 + 2y | 2x + y = 0 -2y ta = = 0, y = 2x + 2xy =0 2x + 2x1-x2) = (2) x/2 - x²) =0 ş